Solve: 2x^4+3x^3+8x^2+15x-10=0
Answer by rothauserc(4718)

2x^4 +3x^3 +8x^2 +15x -10 = 0
:
there will be at most 4 solutions for x
:
use the rational roots theorem since all the coefficients are integers
:
the general form is
:
a(n)x^n +a(n-1)x^(n-1) +.... +a(2)x^2 +a(1)x^1 +a(0) = 0
:
let q = a(n) and p = a(0), then using rational roots theorem we consider factors of p and q
:
q could equal to + or - 1, + or - 2
:
p could be equal to + or - 1, + or - 2, + or - 5, + or - 10
:
p/q = + or - 1, + or - 2, + or - 5, + or - 10, + or - 1/2, + or - 5/2
:
now test these values with the given equation
:
we find that x = 1/2 and x = -2 are rational solutions
:
(x -1/2) * (x +2) = x^2 +1.5x -1 is a polynomial factor of the given equation
:
(2x^4 +3x^3 +8x^2 +15x -10)/(x^2 +1.5x -1) = 2x^2+10
:
2x^2+10 = 0
:
2x^2 = -10
:
x^2 = -5
:
x = i * square root(5) or -i * square root(5)
:
****************************************************
we have two rational roots and 2 complex roots
:
x = 1/2, -2, i * square root(5), -i * square root(5)
****************************************************
: